An unknown salt is added continuously to water until salt is no longer dissolved. Initially, the water weighs 50g. When the mixture reached 65g, salt can no longer be dissolved. What is the saturation concentration of the salt in g/L?


1
Expert's answer
2021-10-06T00:00:39-0400

m(salt) = 65 - 50 = 15 g

saturation concentration = 15 / 50 * 1000 = 300 g/L